OpenStudy (anonymous):

what are the solutions for 3a^2-8=10a i need to solve by factoring or using the principles of 0 products

OpenStudy (anonymous):

3a2-10a-8=0 3a2-12a+2a-8=0 3a(a-4)+2(a-4)=0 (3a+2)(a-4)=0 a=-2/3 or 4

OpenStudy (anonymous):

thank you